# Validity of the Fröhlich model for a mobile impurity in a Bose-Einstein condensate

### v1

@@ I Introduction
The concept of a polaron, a quasi-particle formed through the interaction of an impurity with a medium, is ubiquitous in condensed matter physics [1, 2]. Among the many examples, the Bose polaron stands out by the degree of control over the interactions allowed by cold atomic gases. For this reason, it is seen as a possible platform for the simulation of polaron physics in analogous systems, and in particular allows to access the large-coupling regime [3, 4, 5]. The system also serves a similar role for theory, as a many-body quantum system whose constituents are simple enough that one can understand its complex behavior and effective degrees of freedom from first principles and has thus attracted considerable attention [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. For the same reason, the rigorous justification of the effective description of fluctuations in dilute Bose gases by Bogoliubov phonons is an active area of mathematical research [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], with recent results also considering impurities [34, 35, 36].
@@ III.2 Renormalization of the Fröhlich Hamiltonian
The transformation also creates terms of the form $\sum_{{\bm{k}},{\bm{\ell}}}\hat{V}_{\mathrm{IB}}({\bm{k}})u_{{\bm{k}}+{\bm{\ell}}}v_{\bm{\ell}}b^{\dagger}_{{\bm{k}}+{\bm{\ell}}}b^{\dagger}_{\bm{\ell}}$ from the quadratic interaction term (6d), which are responsible for the logarithmic correction to the energy found in [9]. In our case this is of the lower order $N^{-1/2}\log N$, due to the decay of $v_{k}\sim k^{-2}$ for large $k$ (compare [36, Lem.4.25]).
@@ Acknowledgments
This work was funded by the French National Research Agency (ANR) and the German Research Foundation (DFG) through the projects (MaBoP, ANR-23-CE40-0025; PI 1114/8-1) and CRC-TRR 352 (Project-ID 470903074). J.L. acknowledges additional support through the EUR-EIPHI Graduate School (ANR-17-EURE-0002).

### v2

@@ I Introduction
The concept of a polaron, a quasi-particle formed through the interaction of an impurity with a medium, is ubiquitous in condensed matter physics [1, 2, 3]. Among the many examples, the Bose polaron stands out by the degree of control over the interactions allowed by cold atomic gases. For this reason, it is seen as a possible platform for the simulation of polaron physics in analogous systems, and in particular allows to access the large-coupling regime [4, 5, 6]. The system also serves a similar role for theory, as a many-body quantum system whose constituents are simple enough that one can understand its complex behavior and effective degrees of freedom from first principles and has thus attracted considerable attention [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. For the same reason, the rigorous justification of the effective description of fluctuations in dilute Bose gases by Bogoliubov phonons is an active area of mathematical research [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], with recent results also considering impurities [37, 38, 39].
@@ III.2 Renormalization of the Fröhlich Hamiltonian
The transformation also creates terms of the form $\sum_{{\bm{k}},{\bm{\ell}}}\hat{V}_{\mathrm{IB}}({\bm{k}})u_{{\bm{k}}+{\bm{\ell}}}v_{\bm{\ell}}b^{\dagger}_{{\bm{k}}+{\bm{\ell}}}b^{\dagger}_{\bm{\ell}}$ from the quadratic interaction term (6d), which are responsible for the logarithmic correction to the energy found in [11]. In our case this is of the lower order $N^{-1/2}\log N$, due to the decay of $v_{k}\sim k^{-2}$ for large $k$ (compare [39, Lem.4.25]).
@@ Acknowledgments
This work was funded by the French National Research Agency (ANR) and the German Research Foundation (DFG) through the projects (MaBoP, ANR-23-CE40-0025; PI 1114/8-1) and CRC-TRR 352 (Project-ID 470903074). J.L. acknowledges financial support by the EIPHI Graduate School (ANR-17-EURE-0002) and Bourgogne-Franche-Comté Region through the project SQC.